Circular No. 7835

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ap IN M74 K. S. Kawabata, G. Kosugi, T. Sasaki, Y. Ohyama, N. Kashikawa,

Y. Saito, and M. Iye, National Astronomical Observatory of Japan; and K. Nomoto, University of Tokyo, write: “We carried out sequential optical spectropolarimetry (range 385–830 nm; resolution 0.56 nm and 0.14 nm/pixel) of SN 2002ap (IAUC 7810) on Feb. 9.24, 10.28, 11.26 and 12.26 UT with the Subaru 8.2-m telescope (+ FOCAS). The polarization bump corresponding to the broad absorption feature in the wavelength range 650– 800 nm (cf. IAUC 7820) was still prominent, but the wavelength of the maximum polarization had shifted to 748 nm and the FWHM of the feature had narrowed to 45 nm. A significant variation in the position angle of the polarization with respect to wavelength was observed around this feature. The position angle showed a broad dip of about 15^{o} over the feature at 650–800 nm, superposed on a local (720–770 nm) asymmetric bump of 20^{o} in height that peaked at 760 nm. These findings suggest a complicated asymmetric geometry of the explosion.”

RX J0806.3+1527 G. L. Israel and L. Stella, Astronomical Observatory, Rome; W. Hum-

mel, European Southern Observatory (ESO); and S. Covino and S. Campana, Astronomical Observatory of Brera, on behalf of a larger collaboration, report on spectroscopic and photometric optical observations, obtained with the the ESO Very Large Telescope (on 1999 Nov. 12–19, 2000 Feb. 5, 2001 Jan. 16–24, and Nov. 11) and the Telescopio Nazionale Galileo (on 2001 Jan. 1), of the blue star (α = 8^{h}06^{m}22^{s}.9, δ = +15^{o}27′31′′, equinox 2000.0; V = 21.1) suggested as the optical counterpart of the 321-s x-ray pulsator RX J0806.3+1527 (Israel et al. 1999, A.Ap. 349, L1): “B-, V -, and R-band time-resolved photometry revealed the presence of 15-percent pulsed-fraction modulation at the 321-s x-ray period, confirming the correctness of the identification. Broad (v ∼ 1500 km/s), low equivalent-width (0.1–0.5 nm) emission lines from the He II Pickering series (possibly blended with H Balmer lines) were clearly detected. These findings, together with the period stability and the absence of any additional modulation in the period range 1–300 min, argue in favor of the orbital interpretation of the 321-s pulsations. The most likely scenario is that RX J0806.3+1527 is a double-degenerate system of the AM CVn class. This would make it the binary with the shortest orbital period currently known and one of the best candidates for gravitational-wave detection.”

2002 February 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green